OpenStudy (anonymous):

f: A -> B and g: B -> C are 1-1 functions. Prove that the inverse of g composition f = f inverse composition g inverse on Ran(g composition f)

OpenStudy (anonymous):

\[f: A \rightarrow B \] and \[g: B \rightarrow C \] are one-to-one. Prove that \[[g(f(x))]^{-1} = f^{-1}(g^{-1}(x))\] on \[Ran(g(f(x))\] (someone has to show me how to put all of these on the same line and how to use the little circle for composition D: ) Okay, so here's my brainstorming. I at least want to say what I'm thinking and what I've tried. Okay, so if I say: \[f^{-1}= \left\{ (y,x) \in B \times A | f(x) = y \right\}\] and then I say: \[(g(f(x)) = \left\{ (x,z) | z = g(y) , y = f(x), x \in A\right\}\] Does this mean I can say: \[(g(f(x))^{-1} = \left\{ (z,x) | x = g(y), y= f(z), z \in C \right\}\]??? Assuming that's correct, in order to show equality, I need to show: \[(g(f(x))^{-1} \subseteq f^{-1}(g^{-1}(x))\] AND \[f^{-1}(g^{-1}(x)) \subseteq g(f(x))^{-1}\] correct? One thing I'm definitely not sure on is when it says on Ran(g(f(x)). I guess something like this only holds when we consider the set Ran(g(f(x))? Alright, so I'm assuming I have to say something like let \[z \in (g(f(x))^{-1}\] What that implies, Im not sure. That there exists x = g(f(z)), z in C? I know I'm leaving out the idea of one-to-one, but I just don't know how to have one thing imply another or how to connect the ideas together. All I have written down are a bunch of set definitions with no idea how to piece them together :(

OpenStudy (anonymous):

\circ

OpenStudy (anonymous):

it is always the case that the inverse of \(ab\) is \(b^{-1}a^{-1}\) not matter what the group is

OpenStudy (anonymous):

the only reason we don't notice it is because multiplication is commutative, so it doesn't matter what order you write it, but \[(ab)(b^{-1}a^{-a})=a(bb^{-1})a^{-1}\] as composition is associative, then \[aIa^{-1}\] as \(bb^{-1}=I\) then \[aa^{-1}=I\] and done

OpenStudy (anonymous):

to repeat, has nothing to do with composition as such, true in any associative group

OpenStudy (anonymous):

So, multiplication as an operator in any group is commutative and composition as an operator in any group is associative? And do I not have to show that \[(ab)^{-1} = b^{-1}a^{-1}\] I can just state that?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no i did not say it was commutative, and in fact it is not

OpenStudy (anonymous):

associativity is what is needed to write \[(ab)(b^{-1}a^{-1})=a(bb^{-1})a^{-1}\] that is all

OpenStudy (anonymous):

Oh, you used the word commutative above, maybe you just meant associative then, I'm sorry. So do I need to mention anything about one-to-one or about the Ran(g(f(x)) in any of this?

OpenStudy (anonymous):

in fact if your operation is associative, then the parentheses are not needed, and the one line proof that \((ab)^{-1}=b^{-1}a^{-1}\) is \[abb^{-1}a^{-1}=aea^{-1}=aa^{-1}=e\]

OpenStudy (anonymous):

when i wrote "the only reason we don't notice it is because multiplication is commutative" i meant with numbers for example it doesn't matter \[(ab)^{-1}=a^{-1}b^{-1}\] since multiplication is commutative we can write the inverse in any order we like

OpenStudy (anonymous):

but to show that that works at one point you have to use commutativity not so with \((ab)^{-1}=b^{-1}a^{-1}\) it just works out of the box

OpenStudy (anonymous):

I see. Now I just want to make sure I dont need to make mention of 1-1 or the range portion of the problem.

OpenStudy (anonymous):

maybe you do i think that just makes the composition make sense is all

OpenStudy (anonymous):

Alright. Ill see if I can incorporate it in there. Thanks for you help ^_^

OpenStudy (anonymous):

yw